# LowDiff: Efficient Frequent Checkpointing via Low-Cost Differential for High-Performance Distributed Training Systems

### v1

@@ V-C Design of LowDiff+
Design details. Based on Insight 1 and 2, we design LowDiff+ without gradient compression. The architecture of LowDiff+ is shown in Figure 8, and its detailed algorithm is presented in Algorithm 2. In the training process, the model first performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer as shown in Figure 6 (Line 5). The layer-wise gradient communication is then handled by a communication thread pool $P_{g}$, where the gradient synchronization is assigned to multiple communication threads (Line 15). Once the communication in a thread finishes, the corresponding gradients are forwarded via the reusing queue to the CPU-side checkpointing process (Line 16), while the training process waits for the full gradient $G_{t}$ (i.e., wait for handle set $H_{g}$) (Line 6) to update the GPU model $M^{G}_{t}$ (Line 7).
Figure 8: Workflow of LowDiff without gradient compression.
@@
In the checkpointing process, gradients $G^{n}_{t}$ streamed from the GPU are reused layer by layer (Line 9). The offloading thread pool $P_{s}$ handles multiple layers in parallel (Line 10), snapshotting the layer-wise gradients into CPU memory as they become available (Line 19). We initialize a CPU-resident replica $M^{C}_{t}$ of the model parameters, aligned with the GPU state. After snapshotting (Line 11), reused gradients are applied to $M^{C}_{t}$ to keep it synchronized with the GPU model $M^{G}_{t}$ (Line 12). The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [54]. Since such updates do not involve forward or backward computations, their execution on the CPU introduces negligible overhead. Persistence is performed asynchronously by checkpointing the CPU-maintained model replica (Line 13), rather than the raw gradients. The frequency of persistence is decided similarly to Checkfreq [36], which ensures maximal overlapping between the persist and update in CPU. This design both avoids redundant storage of differential checkpoints and completely decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
@@
When failure occurs, prior work [54, 19] categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures. For hardware failures (e.g., GPU device crashes), the distributed training system must replace the failed machine with a new one. In this case, the in-memory training state is lost, so the recovery process must reload model states from persistent storage checkpoints, which can be time-consuming and increase wasted time. In contrast, software failures (e.g., software bugs) only cause the training process to terminate while leaving the CPU memory state of the checkpointing process intact. In such cases, recovery is achieved by reinitializing the training process and restoring the preserved CPU-resident model state to the GPUs, leveraging in-memory checkpoints. This enables fast recovery under software failures, avoiding expensive checkpoint reloads and serving as a key optimization of LowDiff+.
@@
[…]
VI Implementation
VI-A Implementation of LowDiff
We implement LowDiff atop PyTorch [40, 39] and DeepSpeed [10, 44], consisting of three main modules: reusing queue (§IV-A), batched gradient writing (§IV-B), and optimal configuration (§IV-C). In addition, a parallel recovery module accelerates recovery from a large number of differential checkpoints at high frequency.
@@ VI-A Implementation of LowDiff
Parallel recovery module.Traditional recovery methods for differential checkpointing [15] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (2)), whose the number of merge operations is $n$ with given $n$ differential checkpoints. Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints. Specifically, as shown in figure 9, it merges the earliest differential checkpoint $C_{1}^{D}$ with the full checkpoint $C_{t}^{F}$ while concurrently merging the subsequent differential checkpoints in pairs (i.e., $C_{2}^{D}$ and $C_{3}^{D}$, $C_{4}^{D}$ and $C_{5}^{D}$). This parallelization reduces the number of merge operations between differential checkpoint and full checkpoint from $n$ to $\log n$. In Figure 9, the number of merges between differential checkpoint and full checkpoint is 2 (the serial process has $n=5$ merges). Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
VI-B Implementation of LowDiff+
LowDiff+ is also implemented atop PyTorch [40, 39] and DeepSpeed [10, 44]. Since the checkpointing process is spawned from the training process, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning. For gradient communication, we leverage Python’s concurrent.futures library and employ a ThreadPoolExecutor to assign each layer-wise gradient to a communication thread, ensuring parallel transmission. The snapshotting thread pool is implemented in the same manner, handling the writing of gradients into CPU memory.
To maintain the CPU-side model replica, we configure PyTorch intra-op parallelism via torch.set_num_threads, setting the number of threads to match the available CPU cores in order to maximize parallel efficiency. Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation. Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [36], thereby maximizing checkpointing throughput without interfering with GPU training.
@@ VI-B Implementation of LowDiff+
VII Evaluation
@@ VII-A Experimental Setup
Workloads. We evaluate widely-used DNN models on tasks of image classification and natural language processing (NLP), including ResNet [20], VGG [49], Bert [14], and GPT-2 [41]. The image classification models are evaluated on the CIFAR-100 [25] and ImageNet [12] datasets, and NLP models are evaluated on the SQuAD [43] and WikiText [33] datasets. The detailed configurations are shown in Table II(b). We use Adam [24] as the default optimizer. The training process uses sparsification compression with a common compress ratio $\rho$ of 0.01 [3, 29, 18, 35, 7, 45]. We use default batching sizes and learning rates according to the original papers or standard practices. We also implement the LowDiff system on the VGG-16 model with pipeline parallelism provided in the DeepSpeedExamples [11].
Baselines. We evaluate two state-of-the-art checkpointing strategies: CheckFreq [36] and Gemini [54], as described in §II-B. The experiments follow the default configurations according to their original papers. Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [15] does in recommendation systems.
VII-B Performance Evaluation
@@
Figure 11 shows that the training time of LowDiff+ across all tasks exhibits a slight increase compared to the W/O CKPT baseline, yet it remains significantly lower than that of other checkpointing solutions. Specifically, LowDiff+’s training time increases by 8.2%–10.1% relative to W/O CKPT, which is slightly higher than the increase observed in the setting without gradient compression. This overhead is primarily due to the frequent and large-volume gradient transfers between GPU and CPU memory, which occupy PCIe bandwidth and consequently reduce training performance.
In the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods. For the GPT2-L training task, it reduces the training time by 51.8% compared to Gemini and by 81.7% compared to CheckFreq, demonstrating its superior efficiency in settings where frequent non-compression checkpoints are required.
Figure 12: (Exp. 3) Wasted time under different MTBFs.
Experiment 3 (Wasted time). We evaluate the wasted time of various checkpointing methods in the event of failures. Failures were simulated during training by adhering to a fixed Mean Time Between Failures (MTBF) metric. Using three MTBF values at 0.5, 1, and 2 [19], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S. Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §IV-C. For LowDiff+, we evaluate its wasted time under both software and hardware failures (LowDiff+ (S) and LowDiff+ (H), respectively).
@@
Figure 13: (Exp. 4) Maximum checkpointing frequency.
Experiment 4 (Checkpointing frequency). To evaluate the maximum checkpointing frequency that each method can achieve under bounded training speed (set as 3.5%, as the degraded speed of model training is required to be less than 3.5% in Microsoft’s study [36]), we evaluate the highest checkpointing frequency of different methods. We denote LowDiff+ ’s in-memory checkpointing as LowDiff+ (S) and its persistence as LowDiff+ (P).
@@
Experiment 5 (Recovery time). We evaluate the recovery time of different recovery strategies under different full checkpointing frequency in GPT2-S. We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [39]), Naïve DC, LowDiff with parallel recovery (see §VI), and LowDiff+ from software failures (LowDiff+ (S) for short). Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
@@
Experiment 6 (Checkpointing time reduction and GPU memory reduction). To evaluate the reduction in average checkpointing time achieved by the batched write optimization proposed in §IV-B, we conduct experiments on various models with different batching sizes. Figure 15a shows that the batched write strategy significantly reduces the average time required to write differential checkpoints, with a maximum reduction of $30.9\%$ when the batching size is 20 during GPT2-S training.
The batching strategy effectively accelerates the checkpoint writing process. After tuning the batching frequency with the configuration strategies discussed in §IV-A, the optimization can adapt to different compression ratios and minimize the impact on training performance.
Additionally, we also evaluate the reduction of GPU memory achieved by the batching offloaded in CPU §IV-B. Figure 15b shows that the GPU memory usage will increase by $10\%-12\%$ without offloaded batching, especially for GPT2-L, potentially leading to memory exhaustion. In contrast, the GPU memory usage will return to the original state with the help of offloaded batching.
@@
Compared to Naïve DC, LowDiff further reduces the storage overhead by 90.5%. Because LowDiff utilizes compressed gradients to construct differentials for both model parameters and optimizer parameters (Finding 2 in §III-B), while Naïve DC (Check-N-Run [15]) does not apply sparsification compression to the optimizer parameters within differential checkpoints, resulting in the optimizer occupying the majority of the checkpoint space. The result ensures our motivation that LowDiff can significantly reduce the checkpoint size and thus lower the storage cost.
@@
Experiment 8 (Impact of compression ratio $\rho$). We evaluate the impact of different compression ratios $\rho$ on the checkpoint frequency of GPT2-L and GPT2-S training. The compression ratio $\rho$ varies from 0.1 to 0.001, consistent with the ranges commonly used in prior work [29, 35].
@@ VII-C System Scalability
Experiment 9 (Scaling to frequent failures). We conduct experiments on the performance of LowDiff and LowDiff+ in extreme scenarios characterized by more frequent failures of the training process with MTBF ranging from 0.1 to 5 hours. Using the effective training time ratio metric as defined in Gemini [54], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
Figure 17 demonstrates that LowDiff consistently sustains a higher effective training time ratio across all tested scenarios, especially under environments with frequent failures. Although LowDiff+ achieves slightly lower ratios than LowDiff, it still outperforms the other methods. For example, when the MTBF is set to 0.3 hours, LowDiff reaches an effective training time ratio of 92%, while LowDiff+ achieves 86%, compared to only 81% for Gemini and 76% for Checkfreq. These results highlight the robustness and efficiency of both LowDiff and LowDiff+ in handling frequent failures during training.
Experiment 10 (Scaling to more GPUs). We evaluate the training performance by varying the number of GPUs used in the training process, using the same effective training time ratio metric as in Experiment 7. As the number of GPUs increases, the probability of failures in the cluster also rises [19]. We conduct experiments with 8, 16, 32, and 64 GPUs, respectively, and measure the effective training time ratio of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.

### v2

@@ VI-C Design of LowDiff+
Design details. Based on Insight 1 and 2, we design LowDiff+ under non-compression scenarios. The architecture of LowDiff+ is shown in Figure 9, and its detailed algorithm is presented in Algorithm 2. In the training process, the model first performs backpropagation to calculate gradients $G^{n}_{i,t}$ layer-by-layer as shown in Figure 7 (Line 5). The layer-wise gradient communication is then handled by a communication thread pool $P_{g}$, where the gradient synchronization is assigned to multiple communication threads (Line 15). Once the communication in a thread finishes, the corresponding gradients are forwarded via the reusing queue to the CPU-side checkpointing process (Line 16), while the training process waits for the full gradient $G_{t}$ (i.e., wait for handle set $H_{g}$) (Line 6) to update the GPU model $M^{G}_{t}$ (Line 7).
@@
In the checkpointing process, gradients $G^{n}_{t}$ streamed from the GPU are reused layer by layer (Line 9). The offloading thread pool $P_{s}$ handles multiple layers in parallel (Line 10), snapshotting the layer-wise gradients into CPU memory as they become available (Line 19). This parallel snapshotting allows different layer gradients to be written concurrently to CPU memory, thereby better utilizing the available DRAM bandwidth. We initialize a CPU-resident replica $M^{C}_{t}$ of the model parameters, aligned with the GPU state. After snapshotting (Line 11), reused gradients are applied to $M^{C}_{t}$ to keep it synchronized with the GPU model $M^{G}_{t}$ (Line 12). While gradients are snapshotted layer-by-layer, the CPU-side update (Line 12) is performed once the full gradient is aggregated. This is because the Adam optimizer requires the complete gradient to correctly update its internal states (e.g., momentum and variance). The updated model state in CPU memory serves as an in-memory checkpoint similar to Gemini [wang2023gemini]. Since such updates do not involve forward or backward computations, their execution on the CPU introduces negligible overhead. Persistence is performed asynchronously by checkpointing the CPU-maintained model replica (Line 13), rather than the raw gradients. The frequency of persistence is decided similarly to Checkfreq [mohan2021checkfreq], which ensures maximal overlapping between the persist and update of the model in CPU. This design both avoids redundant storage of differential checkpoints and completely decouples persistence from GPU training, enabling fully asynchronous checkpointing with minimal interference to ongoing training.
Figure 9: Architecture of LowDiff+.
@@
Recovery of LowDiff+. When failure occurs, prior work categorizes failures into two broad types from the perspective of recovery: hardware failures and software failures. For hardware failures (e.g., GPU device crashes), the distributed training system must replace the failed machine with a new one. In this case, the in-memory training state is lost, so the recovery process must reload model states from persistent storage checkpoints, which can be time-consuming and increase wasted time.
In contrast, software failures (e.g., software bugs) only cause the training process to terminate. This design assumes the failure is confined to the training processes (e.g., a CUDA error or NaN loss) and does not corrupt the separate checkpointing process’s memory, thus leaving the CPU memory state of the checkpointing process intact. In such cases, recovery is achieved by reinitializing the training process and restoring the preserved CPU-resident model state to the GPUs, leveraging in-memory checkpoints. This enables fast recovery under software failures, avoiding expensive checkpoint reloads and serving as a key optimization of LowDiff+.
@@
1: GPU model state after the $t$-th iteration $M^{G}_{t}$, gradient of the $n$-th layer on node $i$ $G^{n}_{i,t}$, gradient communication and offload thread pools $P_{g}$ and $P_{s}$, reusing queue $Q$, corresponding handle sets $H_{g}$ and $H_{s}$.
2: function Training Process
3:   for all $t$ do
4:    for all $n$ do
5:      $G^{n}_{i,t}\leftarrow Backward(x^{n}_{t})$
6:      $H_{g}.append(P_{g}.execute(\textsc{Sync},G^{n}_{i,t}))$ 7:                                         $\triangleright$ Sync thread
8:    $G_{t}\leftarrow H_{g}.wait()$ $\triangleright$ Wait for sync
9:    $M^{G}_{t+1}\leftarrow M^{G}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in GPU
10: function Checkpointing Process
11:   while $G^{n}_{t}\leftarrow{\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.get}}$ do $\triangleright$ Zero-copy reuse
12:    $H_{s}.append(P_{s}.execute(\textsc{Snapshot},G^{n}_{t}))$ 13:                                       $\triangleright$ Snapshot thread
14:   $G_{t}\leftarrow H_{s}.wait()$ $\triangleright$ Wait for offload
15:   $M^{C}_{t+1}\leftarrow M^{C}_{t}+Adam(o_{t},G_{t})$ $\triangleright$ Update in CPU
16:   $C^{F}_{t}\leftarrow Save(M^{C}_{t})$ $\triangleright$ Persistence fully
17: function Sync Thread
18:   $G^{n}_{t}\leftarrow Sync(G^{n}_{i,t})$ $\triangleright$ Synchronization
19:   ${\color[rgb]{0.30078125,0.53515625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.30078125,0.53515625,1}\textbf{Q.put}}(G^{n}_{t})$ $\triangleright$ Zero-copy reuse
20:   return $G^{n}_{t}$
21: function Snapshot Thread
22:   $G^{n}_{t}\leftarrow Offload\_to\_cpu(G^{n}_{t})$ $\triangleright$ Snapshot
23:   return $G^{n}_{t}$
VII Implementation
VII-A Implementation of LowDiff
We implement LowDiff atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed], consisting of three main modules: reusing queue (§V-A), batched gradient writing (§V-B), and optimal configuration (§V-C). In addition, a parallel recovery module accelerates recovery from a large number of differential checkpoints at high frequency.
@@ VII-A Implementation of LowDiff
Parallel recovery module.Traditional recovery methods for differential checkpointing [eisenman2022check] use a serial process to merge multiple differential checkpoints with a base full checkpoint (see Equation (6)), whose the number of merge operations is $n$ with given $n$ differential checkpoints. Thus, we implement a parallel recovery module, which parallelizes the merge operations across full and differential checkpoints. Specifically, as shown in figure 10, it merges the earliest differential checkpoint $C_{1}^{D}$ with the full checkpoint $C_{t}^{F}$ while concurrently merging the subsequent differential checkpoints in pairs (i.e., $C_{2}^{D}$ and $C_{3}^{D}$, $C_{4}^{D}$ and $C_{5}^{D}$). This parallelization reduces the number of merge operations between differential checkpoint and full checkpoint from $n$ to $\log n$. In Figure 10, the number of merges between differential checkpoint and full checkpoint is 2 (the serial process has $n=5$ merges). Thus, the parallel recovery accelerates the overall recovery process over the traditional serial one.
@@
VII-B Implementation of LowDiff+
LowDiff+ is also implemented atop PyTorch [pytorch, paszke2019pytorch] and DeepSpeed [deepspeed, rasley2020deepspeed]. Since the checkpointing process is spawned from the training process, the model state in CPU is first initialized by a copy.deepcopy() of the GPU model to ensure consistency at the beginning. For gradient communication, we leverage Python’s concurrent.futures library and employ a ThreadPoolExecutor to assign each layer-wise gradient to a communication thread, ensuring parallel transmission. The snapshotting thread pool is implemented in the same manner, handling the writing of gradients into CPU memory.
To maintain the CPU-side model replica, we configure PyTorch intra-op parallelism via torch.set_num_threads, setting the number of threads to match the available CPU cores in order to maximize parallel efficiency. Importantly, the CPU update step only applies gradients to the model parameters without performing forward or backward propagation, which avoids heavy computation overhead and ensures that using CPU resources does not incur noticeable performance degradation. Finally, the persistence to storage is performed asynchronously, following a style similar to Checkfreq [mohan2021checkfreq], thereby maximizing checkpointing throughput without interfering with GPU training.
VIII Evaluation
@@ VIII-A Experimental Setup
Workloads. We evaluate widely-used DNN models on tasks of image classification and natural language processing (NLP), including ResNet [he2016deep], VGG [simonyan2014very], Bert [devlin2018bert], and GPT-2 [radford2019language]. The image classification models are evaluated on the CIFAR-100 [krizhevsky2009learning] and ImageNet [deng2009imagenet] datasets, and NLP models are evaluated on the SQuAD [SQuAD] and WikiText [merity2016pointer] datasets. The detailed configurations are shown in Table II(b). We use Adam [kingma2014adam] as the default optimizer. The training process uses sparsification compression with a common compress ratio $\rho$ of 0.01 [aji2017sparse, li2022near, fei2021efficient, ming2024adtopk, chen2023bbtopk, renggli2019sparcml]. We use default batching sizes and learning rates according to the original papers or standard practices. We also implement the LowDiff system on the VGG-16 model with pipeline parallelism provided in the DeepSpeedExamples [deepspeedexample].
Baselines. We evaluate two state-of-the-art checkpointing strategies: CheckFreq [mohan2021checkfreq] and Gemini [wang2023gemini], as described in §II-B. The experiments follow the default configurations according to their original papers. Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [eisenman2022check] does in recommendation systems.
VIII-B Performance Evaluation
@@
Figure 12 shows that the training time of LowDiff+ across all tasks exhibits a slight increase compared to the W/O CKPT baseline, yet it remains significantly lower than that of other checkpointing solutions. Specifically, LowDiff+’s training time increases by 7.2%–9.1% relative to W/O CKPT. As expected, this overhead is higher than the 2.4%-3.1% observed for LowDiff in Exp. 1. This increase represents the necessary and expected performance trade-off for handling uncompressed, high-precision gradients, which was the primary motivation for the LowDiff+ design. The overhead is mainly due to frequent and large-volume gradient transfers between GPU and CPU memory, occupying PCIe bandwidth.
Despite this minor overhead, in the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods. For the GPT2-L training task, it reduces the training time by 51.8% compared to Gemini and by 81.7% compared to CheckFreq, demonstrating its superior efficiency in settings where frequent non-compression checkpoints are required.
Experiment 3 (Wasted time). We evaluate the wasted time of various checkpointing methods in the event of failures. Failures were simulated during training by adhering to a fixed Mean Time Between Failures (MTBF) metric. Using three MTBF values at 0.5, 1, and 2 [gupta2024just], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S. Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §V-C. For LowDiff+, we evaluate its wasted time under both software and hardware failures, where the corresponding checkpoints can be accessed throught snapshot and persist ( LowDiff+ (S) and LowDiff+ (P), respectively).
@@
Figure 13: (Exp. 3) Wasted time under different MTBFs.
Experiment 4 (Checkpointing frequency). To evaluate the maximum checkpointing frequency that each method can achieve under bounded training speed (set as 3.5%, as the degraded speed of model training is required to be less than 3.5% in Microsoft’s study [mohan2021checkfreq]), we evaluate the highest checkpointing frequency of different methods. We denote LowDiff+ ’s in-memory checkpointing as LowDiff+ (S) and its persistence as LowDiff+ (P).
@@
Experiment 5 (Recovery time). We evaluate the recovery time of different recovery strategies under different full checkpointing frequency in GPT2-S. We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [paszke2019pytorch]), Naïve DC, LowDiff with parallel recovery (see §VII), and LowDiff+ from software failures (LowDiff+ (S) for short). Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
@@
Experiment 6 (Checkpointing time reduction and GPU memory reduction). To evaluate the reduction in average checkpointing time achieved by the batched write optimization proposed in §V-B, we conduct experiments on various models with different batching sizes. Figure 16a shows that the batched write strategy significantly reduces the average time required to write differential checkpoints, with a maximum reduction of $30.9\%$ when the batching size is 20 during GPT2-S training.
The batching strategy effectively accelerates the checkpoint writing process. After tuning the batching frequency with the configuration strategies discussed in §V-A, the optimization can adapt to different compression ratios and minimize the impact on training performance.
Additionally, we also evaluate the reduction of GPU memory achieved by the batching offloaded in CPU §V-B. Figure 16b shows that the GPU memory usage will increase by $10\%-12\%$ without offloaded batching, especially for GPT2-L, potentially leading to memory exhaustion. In contrast, the GPU memory usage will return to the original state with the help of offloaded batching.
@@
Compared to Naïve DC, LowDiff further reduces the storage overhead by 90.5%. Because LowDiff utilizes compressed gradients to construct differentials for both model parameters and optimizer parameters (Finding 2 in §III-B), while Naïve DC (Check-N-Run [eisenman2022check]) does not apply sparsification compression to the optimizer parameters within differential checkpoints, resulting in the optimizer occupying the majority of the checkpoint space. The result ensures our motivation that LowDiff can significantly reduce the checkpoint size and thus lower the storage cost.
@@
Experiment 8 (Impact of compression ratio $\rho$). We evaluate the impact of different compression ratios $\rho$ on the checkpoint frequency of GPT2-L and GPT2-S training. The compression ratio $\rho$ varies from 0.1 to 0.001, consistent with the ranges commonly used in prior work [li2022near, ming2024adtopk].
@@ VIII-C System Scalability
Experiment 9 (Scaling to frequent failures). We conduct experiments on the performance of LowDiff and LowDiff+ in extreme scenarios characterized by more frequent failures of the training process with MTBF ranging from 0.1 to 5 hours. Using the effective training time ratio metric as defined in Gemini [wang2023gemini], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+. Note that LowDiff+ has two recovery modes: LowDiff+ (S) for software failures (recovering from CPU memory) and LowDiff+ (P) for hardware failures (recovering from persistent storage).
Figure 18 demonstrates that LowDiff+ (S) consistently achieves the highest effective training time ratio across all scenarios, especially under environments with frequent failures. This is because its near-instantaneous recovery from in-memory checkpoints (as shown in Exp. 5) almost completely eliminates recovery overhead, compensating for its slightly higher steady-state overhead compared to LowDiff. LowDiff achieves the second-highest ratio. LowDiff+ (P) also consistently outperforms the other baselines, though its reliance on slower disk-based recovery places it slightly below LowDiff. For example, when the MTBF is set to 0.3 hours, LowDiff+ (S) leads with an effective training time ratio of 94.0%, followed closely by LowDiff at 92%. LowDiff+ (P) achieves 86.8%, which is still significantly higher than Gemini (81%) and Checkfreq (75.9%). These results highlight the robustness of our approaches, with LowDiff+ (S) offering the best performance for common software failures.
Figure 18: (Exp. 9) Training under frequent failures.
Experiment 10 (Scaling to more GPUs). We evaluate the training performance by varying the number of GPUs used in the training process, using the same effective training time ratio metric as in Experiment 7. As the number of GPUs increases, the probability of failures in the cluster also rises [gupta2024just]. We conduct experiments with 8, 16, 32, and 64 GPUs, respectively, and measure the effective training time ratio of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.